40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jan 30 21:29:36 2018, awade, Craig, DailyProgress, PMC, Pre mode cleaner installed in North path 
    Reply  Thu Feb 1 15:25:58 2018, awade, Craig, DailyProgress, PMC, Update working schematic of PMC servo card (older card) D980352-A_AWade20180201Mods.pdf
       Reply  Sat Feb 3 18:08:26 2018, awade, DailyProgress, PMC, A few more modifications to the PMC north board D980352-A_AWade20180202Mods.pdf
          Reply  Fri Jul 6 23:43:54 2018, anchal, DailyProgress, PMC, New modifications to the PMC north board LowPrefFilter.filPMCLowPrefFilterOpt.ipynbU2Buffer.filU2Stability.ipynbD980352-A_AWade20180201Mods_--AGupta20180707Mods.pdf
             Reply  Mon Jul 9 16:15:06 2018, rana, DailyProgress, PMC, New modifications to the PMC north board 
                Reply  Thu Jul 12 15:18:26 2018, anchal, DailyProgress, PMC, New modifications to the PMC north board D980352-AGupta20180712Mods.pdf
                   Reply  Sat Jul 21 14:53:16 2018, awade, DailyProgress, PMC, New modifications to the PMC north board 
                      Reply  Wed Aug 29 12:01:54 2018, anchal, Summary, PMC, Summarization of efforts to fix PMC Servo problems D980352-AGupta20180712Mods.pdfPMC_Servo_LISO_Analysis_Vs_Measured.pdf
                         Reply  Thu Aug 30 12:01:37 2018, rana, Summary, PMC, Summarization of efforts to fix PMC Servo problems 
                         Reply  Mon Sep 24 12:53:26 2018, awade and anchal, Summary, PMC, Summarization of efforts to fix PMC Servo problems: external mixer and LP filter 
Message ID: 2239     Entry time: Mon Sep 24 12:53:26 2018     In reply to: 2228
Author: awade and anchal 
Type: Summary 
Category: PMC 
Subject: Summarization of efforts to fix PMC Servo problems: external mixer and LP filter 

[awade and anchal]

ANCHAL ADD UPDATED TF PLOTS AND PART NUMBERS

For the PMC electronics: this flattening of the TF above 400 Hz was because the on board LP filtering in the RF demodulation stage (LP filter after the mixer) was letting too much RF through in the pass band.  This is a design floor of this particular board that a 1/f^2 filter (made with discrete components) was used instead of a all-in-one 1/f^4 chip.  With too much RF going into the LT1028 there is some kind of saturation problem which causes a leveling off above 400 Hz when taking the TF.

To fix this we put an external minicircuits mixer (model?) followed by a 1/f^4 minicircuits (model?) filter with a corner of 1.9 MHz -- with 50 Ω termination -- into the FP1test point (J3).  At the modulation frequency of 21.5 MHz this should give an attenuation of order -42 dB, much better than before.

[INSERT OLTF]

With the demodulation performed with eternal components the open loop transfer function behaved as expected with the 10 Hz pole setup by the output resistor and PZT working all the way out to very high frequency.  As a result the loop is now stable and working as expected.

 

Quote:

I and Andrew measured transfer functions of each stage of PMC Servo box using Agilent 4395A and HP 41800A probe. Attached are the measured results overplotted with LISO analysis. It seems the PMC Servo box is functioning as it is designed for. Note that in the plots, at higher frequencies there are flat or increasing magnitude regions. These are just because of measurement reaching noise floor. We check independently with sweeps at increased power for higher frequencies and the curves match LISO analysis more or less.
Next, we connected the PZT to the servo box and measured the transfer function with PZT. Again this came out as expected with a good one pole low pass filter behavior as the circuit is designed for. The transfer function (5th plot in the file) keeps going small with a linear slope well up to 10 MHz (measured till this point). So everything checks out fine with he PMC servo box.
But as we lock the servo and measure the closed loop transfer function of the system, we see a strange flat region above 400Hz. Even after the low pass filter with PZT, in a closed loop, it looks like something is actively canceling the pole above 400Hz. But there is nothing else active in the loop. The 6th plot is the measured open loop transfer function (calculated from closed loop transfer function). Note that this flatness is not noise floor or saturation of any kind. We verified this by changing source power and the frequency response at output changes proportionately to this source power change (eg if source power is changed from 5mV to 50mV, the frequency response at the output at 500Hz changes from 4.5mVpkpk to 45mVpkpk), maintaining the flat behavior above 400Hz.
The open loop transfer(G)  function was calculated from closed loop transfer function (C) using this formula:
                                                G = 1 + (1.216)/C
where 1.216 is the measured flat response of the buffer in between the test input (FP2Test) and test output (FP3Test).
So we would like if someone can give some suggestions with this.
Addition: Attaching up-to-date schematic

Quote:

This circuit just doesn't do what it says it should do.  Need to inject waveform at FP1test and probe at each stage.  Then compare against LISO model.  If something is busted we need to know at which point we are getting this extra zero in the response. There is a lot not great about this particular board but it should just have a flat response above 488 Hz.

I don't think AD602 is in the LISO library.  Should be able to add it as some kind of hack op amp with fixed gain, 100nV/rtHz of noise (with 10 Hz corner) and some appropriate current noise with corner of 1 kHz.  Maybe check the AD602 datasheet.

We need to clean this up this week or do something drastic like replacing electronics with minicircuits or removing MC altogether.  For now we need to move onto solving bigger problems like the ISS, thermal stability, PLL readout noise 

Also, on the ISS do you now have a prototype working servo circuit and photodetector?

Quote:

We earlier found that the intended LPF isn't working so we thought of this external LPF idea. So I checked today the PZT input with LCR meter to see if it is in good condition or not. It gave C = 406.4nF, L=60.6 mH and R = 31kOhm. From the spec sheet, the C value looks 20% below the rated value but the spec value has uncertainity of +- 15%, so maybe our PZT is still good.

With the measured values, I calculated again (fitted using LISO) what good value of output resistor would make it closest to 10Hz pole. The value came out to be 37.1k Ohm. I have replaced this output resistor with 39k Ohm now. I'm attaching updated schematic for future reference.

keywords for search: PMC North Driver Board Schematic

Quote:

should probably put back the 30k resistor. It makes a low pass filter with the PZT capacitance (not the cable capacitance). This is to prevent shorting of the HV drive.

 

 

 

 

ELOG V3.1.3-